Calculate the pH of a 0.383 M aqueous solution of nitrous acid (HNO2, Ka = 4.5×10-4) and the equilibrium concentrations of the weak acid and its conjugate base.

pH =
[HNO2 ]equilibrium = M
[NO2- ]equilibrium = M
Answer #1

           HNO2 + H2O H3O+ + NO2-

Initial concentration (M)                                       0.383                         0             0

Change in concentration (M)                                 -X                           X             X

Equilibrium concentration (M)                         (0.383 - X)                      X             X

Given, Ka of HNO2 = 4.5 x 10-4

Now,

Ka = [H3O+][NO2-]/[HNO2]

or, 4.5 x 10-4 = (X)(X)/(0.383 - X)

or, 4.5 x 10-4 = (X)(X)/0.383      [(0.383 - X) 0.383 as X << 0.383]

or, X2 = 0.00017

or, X = 0.013

or, [H3O+] = 0.013

or, -log[H3O+] = -log(0.013)     (taking - log function on both sides of the equation)

or, pH = 1.9

Now, [NO2- ]equilibrium = X = 0.013 M

[HNO2 ]equilibrium = (0.383 - X) = (0.383 - 0.013) M = 0.370 M